LIST OF PARTICIPANTS

Mitsuo Abe Kyoto U, Research Inst for Math Sci, Kitashirakawa Oiwake-cho, Sakyo-ku,

Kyoto 606-01, JAPAN

Ronald Adler San Francisco State U, Physics Dept, 20 Athlene CT, Menlo Park, CA, USA Odylio Aguiar INPE, C.P. 515, SJ Campos, SP 12201-970, BRAZIL Chang-Jun Ahn SLAC, P O Box 4349 Bin 96, Menlo Park, CA 94309, USA Miguel Alcubierre U of Wales, Dept of Physics, College of Cardiff, P.O. Box 913, Cardiff,

Wales CF2 3YB, UK

Gabrielle Allen U of Wales, College of Cardiff, Dept of Physics, P.O. Box 913, Cardiff,

CF2 3YB, UK

Bruce Allen U of Wisconsin, Dept of Physics, P.O. Box 413, Milwaukee, WI, USA Warren G. Anderson U of Alberta, Dept of Physics, Edmonton AB, T6G 2J1, CANADA Paul Anderson Wake Forest U, Physics Dept, P.O. Box 7507, Winston-Salem, NC 27109,

USA

Allen Joel Anderson U of California, Physics Dept, Santa Barbara, CA 93106, USA John D. Anderson Jet Propulsion Lab, 4800 Oak Grove Drive, MS 301-230, Pasadena, CA

91109, USA

Nils Andersson U of Wales, Dept of Physics, College of Cardiff, P.O. Box 913, Cardiff

CF2 3YB, UK

Theocharis Apostolatos Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Marcelo Araujo U of Brasilia, Dept of Matematica, CX Postal 04670, Brasilia D.F., BRAZIL Akito Araya U of Tokyo, 7-3-1 Hongo, Bunkyo, Tokyo 113, JAPAN John Armstrong Jet Propulsion Lab, MS 238-737, 4800 Oak Grove Drive, Pasadena, CA

91109, USA

Jonathan Arons U of California, Astronomy Dept, Berkeley, CA 94720-3411, USA Abhay Ashtekar Penn State U, Gravity Center, 104 Davey Lab, University Park, PA 16802,

USA

Chi Au Hong Kong Polytechnic, Dept of Applied Math, Hung Hom, Kowloon, HONG KONG John Baez U of California, Math Dept, Riverside, CA 92521, USA Roberto Balbinot U of Bologna, Dept of Physics, Via Irnerio, 46, I-40126 Bologna, ITALY Maximo Banados Imperial College, Blackett Lab, Prince Consort Road, London SW7 2BZ,

UK

Michael Bantel U of California, Physics Dept, Irvine, CA 92717, USA Doron Bardas Stanford U, GP-B, Stanford, CA 94305, USA James Bardeen U of Washington, Physics Dept, FM-15, Seattle, WA 98195, USA Mark Barton U of Tokyo, Inst for Cosmic Ray Research, 3-2-1 Midori-cho, Tanashi, Tokyo

203, JAPAN

Andrei Barvinskii U of Alberta, Dept of Physics, Edmonton, Alberta T6G 2J1, CANADA Aroonkumar Beesham U of Zululand, Dept of Applied Maths, Private Bag X1001,

Kwa-Dlangezwa 3886, SOUTH AFRICA

Mark Beilby U of California, Dept of Physics, Irvine, CA 92717, USA Jacob Bekenstein Hebrew U, Racah Inst of Physics, Givat Ram, Jerusalem 91904, ISRAEL Vladimir Belinsky ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Rome, ITALY Baruch Ben-Dor Stanford U, Dept of Aero & Astro, G&C Lab, Stanford, CA 94305, USA John Berberian Stanford U, HEPL, Stanford, CA 94305-8312, USA Beverly K. Berger Oakland U, Physics Dept, Rochester, MI 48309, USA Eric Bergshoeff U of Groningen, Inst for Theor Physics, Nijenborgh 4, 9747Ag Groningen,

THE NETHERLANDS

Alain Bernard ONERA, BP 72, 92322 Chatillon Cedex, FRANCE Orfeu Bertolami CERN, Theory Div, CH-1211 Geneva 23, SWITZERLAND Bruno Bertotti U di Pavia, Dip di Fisica, Via Bassi, 6, I-27100 Pavia, ITALY Massimo Bianchi U di Roma Tor Vergata, Dip di Fisica, Via della Ricerca Scientifica, 1,

I-00133 Rome, ITALY

Andrew Billyard U of Waterloo, Dept of Physics, 200 University Ave W, Waterloo, ON

N2L 3G1, CANADA

Donato Bini ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Rome, ITALY G.S. Bisnovatyi-Kogan Space Research Inst, Profsoyznaya 84/32, 117810 Moscow, RUSSIA David G. Blair U of Western Australia, Dept of Physics, Nedlands, W.A. 6009, AUSTRALIA Luc Blanchet Obs de Paris, 92195 Meudon Cedex, 19 Rue Du Col de Diane, Montigny le

BTX 78180, FRANCE

Elliott Bloom SLAC, MS 98, P.O. Box 4349, Stanford, CA 94309, USA Carles Bona-Garcia U de les Illes Balears, Dep de Fisica, E-07071 Palma de Mallorca, SPAIN Alfio Bonanno U of Alberta, Physics Dept, Edmonton T6G 2J1, CANADA Arvind Borde Tufts U, Inst of Cosmology, Physics Dept, Medford, MA 02155, USA Antonella Borelli ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Roma, ITALY Sukanta Bose U of Wisconsin, Dept of Physics, P.O. Box 413, Milwaukee, WI 53201, USA David Boulware U of Washington, Physics Dept FM-15, Seattle, WA 98195, USA Paul Boynton U of Washington, Dept of Physics, FM-15, Seattle, WA 78175, USA Michael Bradley Umeå U, Dept of Physics, S-90187 Umeå, SWEDEN Howard Brandt Army Research Lab, 2800 Powder Mill Road, Adelphi, MD 20783, USA Steve Brandt NCSA, 605 Springfield Ave, Champaign, IL 61820, USA Hubert Bray Stanford U, Stanford, CA 94305, USA Leo Brewin Minash U, Dept of Math, Clayton, VIX. 3168, AUSTRALIA Alain Brillet CNRS-LAL-VIRGO, Bat 208, Campus, Orsay 91405, FRANCE Roger Brooks MIT, Physics Dept, Cambridge, MA 02139, USA Bernd Bruegmann Max-Planck-Institut, Foehringer Ring 6, D-80805 Mü nchen, GERMANY Saps Buchman Stanford U, Hansen Labs, Stanford, CA 94305, USA S.V. Budnick Astro Space Center, Profsoyznaya 84/32, 117810 Moscow, RUSSIA Daniel Cangemi UCLA, Physics Dept, 405 Hilgard Ave, Los Angeles, CA 90024-1547, USA Giovanni Cantatore U di Trieste, Dip di Fisica, Via A. Valerio, 2, I-34127 Trieste, ITALY Riccardo Capovilla Cinvestav-KPN, Dept of Physics, A. Postal 14-740, 07000 Mexico D.F.,

MEXICO

Paolo Carini Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA Alberto Carlini Nordita, Blegdamsvej 17, 2100 Copenhagen, DENMARK Steven Carlip U of California, Physics Dept, Davis, CA 95616, USA Jaune Carot U de les Illes Balears, Dep de Fisica, E-07071 Palma de Mallorca, SPAIN Gregg Carse CWA, 401 Alberto Way, STE 1, Los Gatos, CA 95032, USA Massimo Cerdonio INFN Natl Lab Legnaro, Via Romea 4, I-35020 Legnaro, ITALY Alberto Chamorro U del Pais Vasco, Dep de Fisica Teorica, Apartado 644, 48080 Bilbao,

Vizcaya, SPAIN

John Chandler Harvard-Smithsonian CfA, MS 63, 60 Garden St, Cambridge, MA 02138,

USA

S. Chandrasekhar U of Chicago, Enrico Fermi Inst, 933 East 56th St, Chicago, IL 60637,

USA

Sujit Chatterjee New Alipore College, Af, Moore Ave, Goverment Housing Estate, Calcutta

700 040, INDIA

Pablo Ag. Chauvet U Autonoma Metropolitana, Av Purisima y Gavilan SN, Mexico D.F.

CP 9300, MEXICO

Anthony Cheng U of Cambridge, DAMTP, Churchill College, Cambridge CB3 ODS, UK Takeshi Chiba Kyoto U, Dept of Physics, Sakyo-ku, Kyoto 606-01, JAPAN Francisco Chinea U Complutense de Madrid, Dep de Fisica Teorica, Fac de Ciencias Fisicas,

28040 Madrid, SPAIN

Matthew Choptuik U of Texas, Center for Relativity, Dept of Physics, Austin, TX

78712-1081, USA

Ernest Chow Manash U, Dept of Math, Wellington Road, Clayton, Victoria 3168,

AUSTRALIA

Demetrios Christodoulou Princeton U, Dept of Math, Princeton, NJ 08540, USA Roberto Cianci U of Genoa, Dept of Math, Via L.B. Alberti, 4, I-16132 Genoa, ITALY David Cline UCLA, Physics and Astronomy, 405 Hillard Ave, Los Angeles, CA, USA Eugenio Coccia U di Roma Tor Vergata, Dip di Fisica, Via della Ricerca Scientifica, 1,

I-00133 Rome, ITALY

Gregory Cook Cornell U, 604 Space Sciences Bldg, Ithaca, NY 14853, USA Fred Cooperstock U of Victoria, Dept of Physics, Victoria, BC V8W 3P6, CANADA Alejandro Corichi Penn State U, Center for Grav Physics, 104 Davey Lab, University Park,

PA 16802, USA

Russell Cosgrove U of California, Dept of Physics, Davis, CA 95616, USA Paolo Cotta-Ramusino U of Milano, Dept of Math, Via Saldini, 50, I-20133 Milano, ITALY Don Cottrell San Diego State U, Dept of Physics, San Diego, CA 92182-0325, USA Ramanth Cowsik Indian Inst of Astrophysics, Koramangala, Bangalore 560 034, INDIA David Craig U of California, Dept of Physics, Santa Barbara, CA 93106-9530, USA Peter D. D’Eath U of Cambridge, DAMPT, Silver St, Cambridge CB3 9EW, UK Thibault Damour Inst des Hautes Etudes Sci, 35 route de Chartres, 91440 Bures sur Yvette,

FRANCE

Karsten Danzmann U of Hannover, Appelstr 2, 30167 Hannover, GERMANY Francesco de Paolis U of Lecce, Dip di Fisica, Via Arnesano CP 193, I-73100 Lecce, ITALY Dan DeBra Stanford, Aero & Astro Dept, MC 4035, Stanford CA 94305, USA Emilio Del Giudice INFN-Milano, Via Celoria 16, I-20133 Milano, ITALY Marek Demianski N Copernicus Astro Center, Bartycka 18, 00716 Warsaw, POLAND Mario C. Diaz Mercyhurst College, Dept of Physics and Astronomy, Erie, PA 16546, USA Honsjorg Dittus ZARM, Center of Applied Space Techonology, AM Fallturm, D-28359

Bremen, GERMANY

Lance Dixon SLAC, Stanford, CA 94305, USA Scott Dodelson Fermilab, NASA, P.O. Box 500, Batavia, IL 60510, USA Tevian Dray Oregon State U, Dept of Math, Corvallis, OR 97331, USA Wolfgang Drechsler Max-Planck-Institut, Föhringer Ring 6, D-80805 Munich, GERMANY Ronald Drever Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Serge Droz U of Alberta, Dept of Physics, Edmonton AB T6G 2J1, CANADA Mark Dubal U of Southampton, Faculty of Math Studies, Southampton S017 1BJ, UK William Duffy Santa Clara Unviersity, Dept of Physics, Santa Clara, CA 95053, USA Frank B. Estabrook Jet Propulsion Lab, Caltech, MS 169-327, Pasadena, CA 91030, USA Francis Everitt Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Edward Fackerell U of Sydney, School of Maths and Stats F07, Sydney NSW 2006,

AUSTRALIA

Viviana Fafone U di Roma Tor Vergata, Dip di Fisica, Via Della Ricerca Scientifica, 1,

I-00133 Rome, ITALY

Fang Li Zhi U of Arizona, Steward Obs, Tucson, AZ 85721, USA Jose Fernandes Monash U, Dept of Math, Wellington Road, Clayton, Victoria 3168,

AUSTRALIA

Valeria Ferrari ICRA, U di Roma, Piazzale Aldo Moro, 2, Roma, 00185, ITALY Simonetta Filippi ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Roma, ITALY David Finkelstein Georgia Inst of Technology, Dept of Physics, Atlanta, GA 30332, USA Eanna Flanagan Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Stephen Fletcher Monash U, Dept of Math, Wellington Road, Clayton, Victoria 3168,

AUSTRALIA

Roberto Floreanini INFN, Dip di Fisica, Strada Costiera, 11, I-34014 Trieste, ITALY William Folkner Jet Propulsion Lab, MS 238-700, 4800 Oak Grove Drive, Pasadena, CA

91009, USA

Lawrence H. Ford Tufts U, Physics Dept, Medford, MA 02155, USA Pierluigi Fortini INFN-Ferrara, Via Paradiso, 12, I-44100 Ferrara, ITALY Roger Foster Naval Research Lab, Code 7210, Washington, DC 20375, USA Carlos Frajuca Louisiana State U, Dept of Physics and Astronomy, Baton Rouge, LA 70803,

USA

Josh Frieman Fermilab, MS 209, P.O. Box 500, Batavia, IL 60510, USA Giorgio Frossati U of Leiden, Galoewater 21, Leiden 2311 VZ, THE NETHERLANDS Francesco Fucito U di Roma Tor Vergata, Dip di Fisica, Via Della Ricerca Scientifica, 1,

I-00173 Roma, ITALY

Yasunori Fujii Nihon Fukushi U, 18011 Tana, Midori-ku, Yokohama, 227, JAPAN Tetsuo Fukui Mukogawa Women’s U, School of Human Life, Ikebiraki-cho, Nishinomiya,

Hyougo 663, JAPAN

Kazumi Fukuma Takuma Natl College of Tech, 551 Kohda, Takuma-cho Mitoyo-gun,

Kagawa 769-11, JAPAN

Wang Tak Fung Hong Kong Polytechnic, Dept of Applied Math, Hung Hom, Kowloon,

HONG KONG

G. Fülöp Chalmers U, Inst Theor Physics, 412 96 Chalmers, Güteborg, SWEDEN Dmitri Gal’tsov Moscow State U, Physics Faculty, 119899 Moscow, RUSSIA Jiang-Gong Gao ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Rome, ITALY Luis J. Garay Imperial College, Theory Group, Blackett Lab, Prince Consort Road, London

SW7 2BZ, UK Juan Garcia-Bellido Stanford U, Physics Dept, Stanford, CA 94305, USA Alberto Garcia-Diaz CIN VEST AV-IPN, Physics Dept, A. Postal 14-740,, C.P. 07000

Mexico D.F., MEXICO

David Garfinkle Oakland U, Dept of Physics, Rochester, MI 48309, USA Charles A. Garwin 1317A Franklin St, Santa Monica, CA 90404, USA Jerome Gauntlett U of Chicago, Enrico Fermi Inst, 5640 South Ellis Ave, Chicago, IL 60657,

USA

Claude Gauthier U de Moncton, Dept de Math, Moncton, NB E1A 3E9, CANADA Kenny Geng Louisiana State U, Dept of Physics, Baton Rouge, LA 70803, USA Ulrich Gerlach Ohio State U, Dept of Math, 231 W 18th Ave, Columbus, OH 43210, USA Giacomo Giampieri Jet Propulsion Lab, Caltech, 4800 Oak Grove Drive, Pasadena, CA

91109, USA

Mauro Giavalisco Space Telescope Science Inst, 3700 San Martin Drive, Baltimore, MD

21218, USA

Adalberto Giazotto INFN-Pisa, U di Pisa, Via Vecchia Livornese 582/A, I-56010 S. Piero a

Grado (PISA), ITALY

Bernard Gilroy Stanford U, Physics Dept, MC 4060, Stanford, CA 94305, USA Edward Glass U of Windsor, Physics Dept, Windsor, Ontario N9B 3P4, CANADA Biping Gong Gaungzhou Medical College, Dept of Physics, Guangzhou 510182, CHINA Gabriela Gonzalez Syracuse U, Dept of Physics, 201 Physics Bldg, Syracuse, NY 13244-1130,

USA

Luis Manuel Gonzalez-Romero U Complutense de Madrid, Dept Fisica Teorica, Fac de

Ciencias Fisica, 28040 Madrid, SPAIN

Daniel Gottlieb Purdue U, Math Dept, 1160 Camelback Blvd, West Lafayette, IN 47907,

USA

Stefan Gottlöber Astrophysical Inst, An der Sternwarte 16, D-14482 Potsdam, GERMANY Pierre Gravel Natl Defence, CMR St. Jean (Richelain), Quebec J0J 1R0, CANADA Jorge Griego INFN, Tistan Narvaja 1674, Montevideo 11200, URUGUAY Jerry Griffiths Loughborough U, Dept of Math Sciences, U of Technology, Loughborough,

Leics, LE11 3TU, UK

Frank Gronwalch U of Texas, 2103 Nueres, Austin, TX 78705, USA Jens Gundlach U of Washington, Nuclear Physics Lab GL-10, Seattle, WA 98195, USA Herbert Gursky Naval Research Lab, Space Science Div, 4555 Overlook Ave SW,

Washington, DC 20375-5320, USA

Eric Gustafson Stanford U, Ginzton Lab, Box 32, Stanford, CA 94305, USA Jemal Guven U Nacional Autonoma, A. Postal 70-543, Mexico, D.F., MEXICO Jason Gwo Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA Roger Halling Umeå U, Dept of Physics, S-90187 Umeå, SWEDEN Leopold Halpern Florida State U, Dept of Physics, Tallahassee, FL 32306-3016, USA William O. Hamilton Louisiana State U, Dept of Physics, Baton Rouge, LA 70803-4001, USA Richard Hammond North Dakota State U, Physics Dept, Fargo, ND 58105, USA Hugh Harleston UNAM, Inst de Ciencias Nucleares, C.U. A. Postal 70-543, 04510 Mexico,

D.F., MEXICO

William Harper U of Western Ontario, Philosophy Dept, London, Ontario, CANADA Tina Harriott Mount Saint Vincent U, 166 Bedford Highway, Halifax, Nova Scotia,

CANADA

Steven Harris St Louis U, Dept of Math, St Louis, MO 63103, USA Gregory Harry U of Maryland, 410 Ridge Road #6, Greenbelt, MD 20770, USA James Hartle U of California, Dept of Physics, Santa Barbara, CA 93106, USA Gordon Haupt Stanford U, GP-B, Hansen Labs, Stanford, CA 94305-8312, USA Sean Hayward Kyoto U, Dept of Physics, Kyoto 606-01, JAPAN Yaron Hefetz MIT, Room 20B-145, Cambridge, MA 02139, USA Michael Heifetz Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Ron Hellings Jet Propulsion Lab, 4800 Oak Grove Drive, Pasadena, CA 91109-8099, USA Thomas Helliwell Harvey Mudd College, Dept of Physics, Claremont, CA 91711, USA Rhett Herman Montana State U, Dept of Physics, 9631 Cougar Drive, Bozeman, MT 59715,

USA

Toshinari Hirai Waseda U, Dept of Physics, Okubo 3-4-1 Shinjuku, Tokyo, 169, JAPAN Jun Ichi Hirao Inst of Space and Astro Sci, 3-1-1 Yoshinodai, Sagamihara, Kanagawa 229,

JAPAN

Kouichi Hirotani Kyoto U, Dept of Physics, Oiwake-cho, Kitashirakawa, Sakyo-ku, Kyoto

606-01, JAPAN

William Hiscock Montana State U, Physics Dept, Bozeman, MT 59717, USA David Hobill U of Calgary, Dept of Physics, Calgary, Alberta T2N 1N4, CANADA James H. Horne Yale U, Dept of Physics, 217 Prospect St, New Haven, CT 06520-8120, USA Gerald Horwitz Hebrew U, Jerusalem, ISRAEL James Hough U of Glasgow, Dept of Physics, University Ave, Glasgow G12 8QQ, UK Yinglei Huang Syracuse U, Dept of Physics, 201 Physics Bldg, Syracuse, NY 13244-1130,

USA

Scott Hughes Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Sheichi Ichinsoe U of Shizuoka, Yada 52-1, Shizuoka 422, JAPAN Gabriele Ingrosso U di Lecce, Dip di Fisica, Via Arnesano C.P. 193, I-73100 Lecce, ITALY James Ipser U of Florida, Dept of Physics, Williamson Hall, Gainesville, FL 32611, USA Hideki Ishihara Kyoto U, Dept of Fundamental Sciences, FIHS, Yoshida Nihonmastu-cho,

Kyoto, 606-01, JAPAN

V.D. Ivashchuk Surface and Vacuum Research, 8 Kravchenko St, 117331 Moscow, RUSSIA Junichi Iwasaki U of Pittsburgh, Physics Dept, Pittsburgh, PA 15260, USA Mark W. Jacobs Stanford U, Physics Dept, Stanford, CA 94305, USA Yusuf Jafry SSD, ESTEC, EXA, P.O. Box 299, 2200 AG Noordwijk, THE NETHERLANDS Robert Jantzen Villanova U, Dept of Math Sciences, Villanova, PA 19085, USA Markus Jarasch Portland State U, C/O Adam Coutu, 2449 SE Taylor, Portland, OR 97214,

USA

Philippe Jetzer U of Zürich, Inst for Theor Physics, Winterthurerstr 190, CH-8057 Zürich,

SWITZERLAND

Warren W. Johnson Louisiana State U, Dept of Physics, Baton Rouge, LA 70803-4001, USA Christine Jones Harvard-Smithsonian CfA, 60 Garden St, Cambridge MA 02138, USA Ramesh Joshi 33076 Marsh Hawk CT, Union City, CA 94587, USA Yongduk Kim Sogang U, Physics Dept, 1-1 Shinsoodong, Mapoku, KOREA Osamu Kaburaki Tohoku U, Astronomical Inst, Aramaki, Aoba-ku, Sendai 980-77, JAPAN Dimitri Kalligas Stanford U, Hansen Labs, GP-B, Palo Alto, CA 94305, USA Renata Kallosh Stanford U, Physics Dept, Stanford, CA 94305, USA Henry Kandrup U of Florida, Dept of Physics, Gainesville, FL 32611, USA Gungwon Kang U of Maryland, Dept of Physics, College Park, MD 20742, USA Arie Kapulkin U of Texas, Center for Relativity, RLM 9.206, 26th and Speedway, Austin,

TX 78712-1081, USA

Jeremy Kasdin Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA David Kastor U of Massachusetts, Physics Dept, Amherst, MA 01003, USA Louis H. Kauffman U of Illinois, Math Dept, 851 South Morgan St, Chicago, Il 60607-7045,

USA

Masoud Kavian U of Waterloo, Dept of Applied Math, Waterloo, ON N2L 3G1, CANADA Nobuki Kawashima Inst of Space and Astronautics, 3-1-1 Yoshinodai, Sagamihara,

Kanagawa 229, JAPAN

Demosthones Kazanas NASA, Goddard Space Flight Center, Code 665, Greenbelt, MD

20771, USA

Mac Keiser Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Daniel Kennefick Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Thomas Kephart Vanderbilt U, Dept of Physics and Astronomy, Nashville, TN 37235, USA David Kerrick Philadelphia College of Pharmacy and Science, 600 S 43rd St, Philadelphia,

PA 19104-4495, USA

Sung-Won Kim EWHA Womans U, Dept of Science Education, Seoul 120-750, KOREA Sang Pyo Kim Kunsan Natl U, Dept of Physics, Kunsan 573-360, KOREA A. A. Kirillov Inst of Applied Math, Uljanova St 10, 603005 Nizhny Novgorod, RUSSIA Klaus Kirsten U de Barcelona, D’ECM, Kiagonal 647, 08028 Barcelona, SPAIN Christian Klein Max-Planck-Inst, Arbeits Gruppe Gravitation, Max-Wien-Platz 1, D-07743

Jena, GERMANY

Wilhelm Kley Friedrich-Schiller U, MPG-AG Gravitation, Max-Wien-Platz 1, D-07743 Jena,

GERMANY

Shiho Kobayashi Tokyo Inst of Technology, Dept of Physics, Hasaya Lab, Oh-Okayama

2-12-1, Meguro-ku, Tokyo 152, JAPAN

Tatsuhiko Koike Tokyo Inst of Technology, Dept of Physics, Oh-Okayama 2-12-1,

Meguro-ku, Tokyo, 152, JAPAN

Yasufuni Kojima Tokyo Metropolitan U, Dept of Physics, Hachioji, Tokyo 192-03, JAPAN Edward Kolb Fermilab, Box 500, Batavia, IL 60510, USA B.V. Komberg Astro Space Center, Profsoyznaya 84/32, 117810 Moscow, RUSSIA D.A. Kompaneets Astro Space Center, Profsoyznaya 84/32, 117810 Moscow, RUSSIA Deborah Konkowski U.S. Naval Academy, Dept of Physics, Annapolis, MD 21402, USA Yasuhiro Koyama Kashima Space Research Center/CRL, 893-1, Hirai, Kashma, Ibaraki,

314, JAPAN

Dietrich Kramer Friedrich-Schiller U, Theor-Phys Inst, Max-Wien-Platz 1, D-07743 Jena,

GERMANY

Nadathur Krishnan U of California, Physics Dept, Irvine, CA 92715, USA Pavel Krtous U of Alberta, 412 Avadh Bhatia Physics Lab, Edmonton, T6G 2J1 Alberta,

CANADA

Lev Kufman U of Hawaii, Inst for Astronomy, 2680 Woodlawn Dr, Honolulu, HI 96822, USA Hannu Kurki-Suonio U of Helsinki, Research Inst for Theor Physics, P.O. Box 9, 00014

Helsinki, FINLAND

Kazuaki Kuroda Inst for Cosmic Ray Research, 3-2-1, Midori-cho, Tanashi, Tokyo 188,

JAPAN

Claus Laemmerzahl U of Konstanz, Postfach 5560, M 674, D-78434 Konstanz, GERMANY Yu Yiu Lam Hong Kong Polytechnic, Dept of Applied Math, Hung Hom, Kowloon,

HONG KONG

Benjamin Lange Stanford U, 1922 Page St, San Francisco, CA 94117, USA Kristine Larsen Central Connecticut State, Physics & Earth Sciences Dept, 1615 Stanley St,

New Britain, CT 06053, USA

Shane L. Larson Montana State U, 200 Quan, Belgrade, MT 59714, USA Eunice Lau Jet Propulsion Lab, MS 301/150, 4800 Oak Grove Drive, Pasadena, CA

91109-8099, USA

Dustin Laurence Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Philippe Laurent DAPNIA/SAP, CEN Saclay, 91191 Gif sur Yvette Cedex, FRANCE George Lavrelashvili U of Zürich, Inst for Theor Physics, Winterthurerstrasse 190, CH-8057

Zürich, SWITZERLAND

Daniel Lebach Harvard-Smithsonian CfA, MS 10, 60 Garden St, Cambridge, MA 02130, USA Wa Tsan Lee Hong Kong Polytechnic, Applied Maths, Hung Hom, Kowloon, HONG KONG Dana Lehr Stanford U, Physics Dept, Stanford, CA 94305, USA José Lemos Obs Nacional, Rua General Jose Cristino 77, 20921 Rio Janeiro, BRAZIL Ron Lenk 100 Pasito Terrace #111, Sunnyvale, CA 94086, USA Melvin Lewis 18 Walnut St, Oakland, NJ 07436, USA Lee Lindblom Montana State U, Physics Dept, Bozeman, MT 59717, USA Andrei Linde Stanford U, Physics Dept, Stanford, CA 94305, USA Nick Linthorne U of Western Australia, Dept of Physics, Nedlands, WA 6009, AUSTRALIA John Lipa Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA Chong-yin Lo Applied & Pure Research Inst, 315 Whytegate CT, Lake Forest, IL 60045,

USA

James Lockhart Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA Jennifer Logan U of Glasgow, Dept of Physics, University Ave, Glasgow G12 8QQ, UK Renate Loll INFN-Firenze, Largo E. Fermi, 2, I-50125 Firenze, ITALY Daniel Loranz 407 West Harrison, Bozeman, MT 59705, USA Darren B. Lortan U of Natal, Dept of Math, King George V Ave, Durban 4001, SOUTH

AFRICA

David Lowe U of California, Dept of Physics, Santa Barbara, CA 93106, USA Hugh Luckock U of Sydney, School of Math, Sydney NSW 2006, AUSTRALIA Anthony W.C. Lun Monash U, Dept of Math, Clayton, VIC 3168, AUSTRALIA Jun Luo Huazhong Univ of Science & Technology, Wuham 430074, CHINA Andrew Lyne U of Manchester, Jodrell Bank, Macclesfield SK11 9DL, UK Harald Lück Max-Planck-Institut, Aupenstelle Hannover, Appelstr. 2, 30167 Hannover,

GERMANY

Grant Mathews Lawrence Livermore Natl Lab, L-294, Livermore, CA, USA Ellchi Mizuno Inst of Space & Astronautical Sciences, 3-1-1 YOSHINODAI, SAGAMIHARA

229, JAPAN Takashi Mishima Nihon U, Lab of Physics, 7-24-1 Narashinodai, Funabashi, Chiba, 274,

JAPAN

Duccio Macchetto Space Telescope Science Inst, 3700 San Martin Drive, Baltimore, MD

21218, USA

Alfredo Macias U A. Metropolitana, Physics Dept, P.O. Box 55-534, 09340 Mexico D.F.,

MEXICO

Kei-ichi Maeda Waseda U, Dept of Physics, Okubo 3-4-1 Shinjuku, Tokyo 169, JAPAN Michele Maggiore INFN, Dip di Fisica, Piazza Torricelli, 2, I-56100 Pisa, ITALY Seth Major Penn State U, Center for Grav Physics, 104 Davey Lab, University Park, CA

16802, USA

Jerome Malenfant Physical Review Letters, 1 Research Road, Box 1000, Ridge, NY 11961,

USA

Catherine Man CNRS, VIRGO-LAL, Bldg 208 Campus, Orsay 91405, FRANCE Gianpiero Mangano U of Naples, Physics Dept, Mostra D’Oltremare Pad 19, I-80125 Napoli,

ITALY

Philip Mannheim U of Connecticut, Dept of Physics, Storrs, CT 06269-3046, USA Mattias Marklund Umeå U, Dept of Physics, S-90187 Umeå, SWEDEN Donald Marolf Penn State U, Dept of Physics, 104 Davey Lab, University Park, PA 16802,

USA

Maurizio Martellini U di Milano, Physics Dept, Via Celoria, 16, I-20133 Milano, ITALY Erik Martinez Theor Physics Inst, U of Alberta, Edmonton, AB T6G 2J1, CANADA Alessandro Massarotti Harvard-Smithsonian CfA, MS 51, 60 Garden St, Cambridge, MA

02138, USA

Joan Masso NCSA, 605 Springfield Ave, Champaign, IL 61820, USA John Mather Princeton U, Dept of Math, Princeton, NJ 08540, USA Tonatiuh Matos Chassin CINVESTAV, Dept of Physics, A. Postal 14-740, Mexico City

07000 DF, MEXICO

David Matravers U of Portsmouth, Schl of Math Studies, Portsmouth, Hants. P01 2EG, UK George Matsas Inst de Fisica Teorica, Rua Pamplona 145, Sao Paulo, SP 01405-900, BRAZIL Petrisor Mazilu Technische Hochschule Darmstadt, Petersen Str. 30, D-64287 Darmstadt,

GERMANY

Giovanni Mazzitelli INFN, P.O. Box 13, I-00044 Frascati, ITALY David McClelland Australian Natl U, Dept of Physics, Canberra 0200, AUSTRALIA John McCuan Stanford U, Math Dept, Stanford, CA 94025, USA Raymond McLenaghan U of Waterloo, Dept of Applied Math, 200 University Ave W,

Waterloo, ON N2L 3G1, CANADA

Renata Mele D’Agnese U of Milan, Dept of Physics, Via Celoria, 16, I-20133 Milano, ITALY Maria Teresa Menna ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Rome, ITALY Pietro Menotti U di Pisa, Dip di Fisica, Piazza Torricelli, 2, I-56100 Pisa, ITALY Marco Merafina ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Roma, ITALY John Mester Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA Peter Michelson Stanford U, Physics Dept, Stanford CA 94305, USA Gennaro Miele U of Naples, Physics Dept, Mostra D’Oltremare Pad 19, I-80125 Napoli,

ITALY

Warner Miller Los Alamos Natl Lab, T-6 Group, MS B-288, Los Alamos, NM, USA Nikolai V. Mitskievich Dep de Fisica, U de Guadalajara, A. Postal 1-2011, C.P. 44100

Guadalajara, Jalisco, MEXICO

Masayoshi Mizouchi Okayama U of Science, Dept of Applied Physics, Ridai-cho 1-1,

Okayama-chi, Okayama 700, JAPAN

Yuko Mochizuki Rikkyo U, Dept of Physics, Nishi-Ikebukuro, Toshima-ku, Tokyo 171,

JAPAN

Marc Moniez CNRS, Lab de l’Accelerateur Lineaire, Bat 200, 91405 Orsay Cedex, FRANCE Paolo Vargas Moniz U of Cambridge, DAMTP, Silver St, Cambridge, CB3 9EW, UK Giovanni Montani ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Roma, ITALY Marcos Montes Stanford U, Physics Dept, Stanford, CA 94305-4060, USA Michael Moore U of Washington, Dept of Physics, FM-15, Seattle, WA 98195, USA Hugo A. Morales-Tecotl U Autonoma Metropolitana, Av. Michoacan y Purisima, Col.

Vicentina, P.O. Box 55-534, Mexico D.F. CP 09340, MEXICO

René Frédéric Morand L.A.P.P., BP 110, 74940 Annecy le Vievx, Cedex, FRANCE Peter Morse U of California, Dept of Physics, Santa Barbara, CA 93117, USA Benoit Mours L.A.P.P., BP 110, 74940 Annecy le Vieux, Cedex, FRANCE Barry Muhlfelder Stanford U, Physics Dept, GP-B, Stanford, CA 94305, USA Ivan Muzinich U of Southern California, Scaver Science Center, Dept of Physics, Los

Angeles, CA 90089-0484, USA

Robert Myers McGill U, Dept of Physics, 3600 University Ave, Montreal, Quebec H3A 2T8,

CANADA

Jürgen Müller Technische Univ München, Forschungseinrightung Satellifengeodisie, D-80290

München, GERMANY

Kouji Nakamura Nagoya U, Dept of Physics, Chikusa-ku, Nagoya 464-01, JAPAN Takashi Nakamura Kyoto U, Dept of Physics, Sakyo-ku, Kyoto, 606-01, JAPAN Ken-ichi Nakao Kyoto U, Dept of Physics, Sakyo-ku, Kyoto, 606-01, JAPAN Shin-Ichi Nakariki Okayama U of Science, Dept of Applied Physics, Ridai-cho 1-1, Okayama

700, JAPAN

Yasusada Nambu Enrico Fermi Inst, U of Chicago, 560 Ellis Ave, Chicago, IL 60637, USA N.B. Narojnyi Moscow Engineering Physical Inst, Kashirskoe Shosse 31, 115409 Moscow,

RUSSIA

Yuval Ne’eman U of Texas, Dept of Physics, Austin, TX 78712, USA James M. Nester Natl Central U, Dept of Physics, Chung-Li 32054, TAIWAN Gernot Neugebauer Max-Planck-Gesellschaft, AG Gravitationstheorie, Max-Wien-Platz 1,

D-07743 Jena, GERMANY

David Newell U of Colorado, JILA, Campus Box 440, Boulder, CO 80309, USA Skip Newhall Jet Propulsion Lab, 4800 Oak Grove Drive, Pasadena, CA 91109-8099, USA Riley Newman UC Irvine, Dept of Physics, Irvine, CA 92717, USA Wei-Tou Ni Natl Tsing Hua U, Dept of Physics, Hsinchu, Taiwan 30043, TAIWAN David Nicholson U of Wales, College of Cardiff, The Parade, Cardiff, Wales CF2 3YB, UK Ulf Nilsson Stockholm U, Dept of Physics, Box 6730, S-11385 Stockholm, SWEDEN Shogo Nishida U of Tokyo, Dept of Astronomy, Komaba, Meguro-ku, Tokyo 153, JAPAN Hyerim Noh Korea Astronomy Obs, San 36-1, Huaam-dong, Yousung-gu, Daejon 305-348,

KOREA Hans-Peter Nollert U Tuebingen, Theor Astrophysik, Auf der Morgenstelle 10C, 27076

Tuebingen, GERMANY

Ken Nordtvedt Montana State U, Northwest Analysis, 118 Sourdough Ridge, Bozeman MT

59715, USA

Pierre H. Noyes SLAC, P.O. Box 4349, Stanford, CA 94309, USA Octavio Obregon U de Guanajuato, Inst Fisica, Loma del Bosque S/N/, Fracc. Lomas del

Campestre, Leon, Guanajuato 37150, MEXICO

Masatake Ohashi Natl Astronomical Obs, 2-21-1 Osawa, Mitaka, Tokyo 181, JAPAN Teruya Ohtani Kansai Gaidai U, 16-1 Kitakatahoko-cho, Hirakata, Osaka 573, JAPAN Ken Olum M.I.T., 156 Massapoag Ave, Sharon, MA 02067, USA Tomas Ort’n Queen Mary and Westfield College, Dept of Applied Maths, Mile End Road,

London E1 4NS, UK

James Overduin U of Victoria, P.O. Box 3055, Victoria, BC V8W 3P6, CANADA Ben Owen Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Don N. Page U of Alberta, 412 Physics Lab, Edmonton, Alberta T6G 2J1, CANADA Ho Jung Paik U of Maryland, Dept of Physics, College Park MD 20742, USA Giovanni Vittorio Pallottino U di Roma, Dept of Physics, Piazzale Moro, 2, I-00185 Rome,

ITALY

Wolfgang Panofsky Stanford U, SLAC, Stanford, CA 94309, USA Philippos Papadopoulos U of Pittsburgh, Dept of Physics, 100 Allen Hall, Pittsburgh, PA

15260, USA

Giorgio Papini U of Regina, Physics Dept, Regina, Sask. S4S OA2, CANADA Leonard Parker U of Wisconsin, Physics Dept, Milwaukee, WI 53201, USA Richard Parmley Lockheed, 3251 Hanover St, Bldg 253, Palo Alto, CA 94304, USA Diego Pavon U Barcelona, Dept of Physics, 08193 Bellaterra, SPAIN Amanda Peet Stanford U, Physics Dept, Stanford, CA 94305-4060, USA Ue-Li Pen Princeton U, Dept of Physics, Princeton, NJ 08540, USA Hong Peng U of Western Australia, Dept of Physics, Nedlands, WA 6009, AUSTRALIA Qiu-he Peng Nanjing U, Dept of Astronomy, Nanjing 210008, CHINA Volker Perlick TU Berlin, Inst Theor Physics, Sekr. PN 7-1, D-10623 Berlin, GERMANY Alain Perret C.N.E.S., FRANCE Phil Perry U of Victoria, Box 3055, Victoria, BC V8W 3P6, CANADA Glenn Peterson U of Texas, Centre for Space Research, 2901 NIH 35-STE 300, Austin, TX

78722, USA

Arlie Petters Princeton U, Dept of Math, Princeton, NJ 08544, USA Luis Octavio Pimentel U Autonoma Metropolitana, P.O. Box 55-534, CP. 9340 Mexico D.F.,

MEXICO

Innocenzo M. Pinto U of Salerno, Via Ponte dou Melillo, I-84044 Salerno, ITALY Fabrizio Pinto Boise State U, Dept of Physics SN-318, 1910 U Drive, Boise, ID 83725, USA Nelson Pinto Neto UPMC, Lab de Gravitation, Tour 22-12, 4, Roite 142, 4 Place Jussieu,

75252 Paris Cedex 05, FRANCE

William Pitter U del Zulia, Dep de Fisica, Maracaibo, Zulia, VENEZULA Guido Pizzella U di Roma, Dept of Physics, Piazzale Aldo Moro, 2, I-00185 Rome, ITALY Jerzy Plebanski Centro de Investigacion y de Estudies Avamzados, A. Postal 14-740, Mexico

07000, D.F., MEXICO Eric Poisson Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Stephen Poletti U of Adelaide, Physics Dept, Adelaide, S. Australia 5001, AUSTRALIA Alexandre Polnarev Queen Mary and Westfield College, Dept of Appl Maths, Mile End

Road, London E1 4NS, UK

Daniel Prager Monash U, Dept of Math, Wellington Road, Clayton, Victoria, 3168,

AUSTRALIA

Giuliano Preparata INFN, via Celoria, 16, I-20133 Milano, ITALY Jorge Pullin Penn State U, Dept of Physics, 104 Davey Lab, University Park, PA 16802, USA Hisham A. Qayum U of California, Dept of Physics, Davis, CA 95616, USA Letao Qin Montana State U, Physics Dept, Bozeman, MT 59717, USA Hernando Quevedo UNAM, A. Postal 70-543, Mexico D.F., MEXICO Oleg Puzyrko Stanford U, Physics Dept, Stanford, CA 94305, USA Mohan Rajagopal Stanford U, Physics Dept, Stanford, CA 94305, USA Malik Rakhmanov Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Jim Rankin Rankin Consulting, 228 W 71st St, Apt. 4J, New York, NY 10023, USA Michael Ratner Harvard-Smithsonian CfA, MS 42, 60 Garden St, Cambridge, MA 02138,

USA

Alpan Raval U of Maryland, Dept of Physics, College Park, MD 20742, USA Mark Redar San Francisco State U, 1600 Holloway Ave, San Francisco, CA 94117, USA Rüdeger Reinhard ESTEC, Space Science Dept, Postbus 299, 2200 AG Noordwijk, THE

NETHERLANDS

Michael Reisenberger Washington U, Dept of Physics, St Louis, MO 63130, USA Paul Renteln California State U, Dept of Physics, 5500 University Pkwy, San Bernardino,

CA 92407-7005, USA

Martin Reuter DESY-T, Notkestrasse 85, D-22603 Hamburg, GERMANY Jean-Paul Richard U of Maryland, Dept of Physics, College Park, MD 20742, USA Sam Richman U of Colorado, JILA, Campus Box 440, Boulder, CO 80309-0440, USA Rogers Ritter U of Virginia, Dept of Physics, McCormick Road, Charlottesville, VA 22901,

USA

David Robertson Glasgow U, Dept of Physics, University Ave, Glasgow G12 8QQ, UK Thomas Roman Tufts U, Dept of Physics, Medford, MA 02155, USA Roger Romani Stanford U, Physics Dept, Stanford, CA 94305-4060, USA M.M. Romanova Space Research Inst, Profsoyznaya 84/32, 117810 Moscow, RUSSIA Marcos Rosenbaum UNAM, A. Postal 70-543, 04510 Mexico, D.F., MEXICO Kjell Rosquist Stockholm U, Dept of Physics, Box 6730, S-11385 Stockholm, SWEDEN Graham Ross Lockheed Missiles & Space Co., 0/92-05 B/253, 3251 Hanover St, Palo Alto,

CA 94304, USA

Haret Rosu IFUG, Loma Del Bosque #131, A. Postal E-143, Leon, Guanajuato, Gto.,

MEXICO

Sheila Rowan U of Glasgow, Dept Physics and Astronomy, Glasgow G12 8QQ, Scotland, UK Albrecht Rüdiger Max-Planck-Institut, Hans-Kopfermann-Str 1, D-85748 Garching,

GERMANY

Remo Ruffini ICRA, U di Roma, Piazzale Aldo Moro, 2, I-00185 Rome, ITALY Michael Ryan UNAM, A. Postal 70-543, Mexico D.F. 04510, MEXICO Fintan Ryan Caltech, Theor Astrophysics, 130-33, Pasadena, CA 91125, USA Mario Spinetti INFN-Frascati, Via Fermi, 13, I-00044 Frascati, ITALY Masa-aki Sakagami Fukui U, Fac of Education, Bunkyo 3-9-1, 910 Fukui, JAPAN Nobuyuki Sakai Waseda U, Dept of Physics, 3-2-1 Okubo, Shinjuku-ku, Tokyo 169, JAPAN Patricio Salgado U de Concepcion, Fac de Fisicas, Casilla 4009, Concepcion, CHILE Roberto Salgado Syracuse U, 201 Physics Bldg, Syracuse, NY 13244-1130, USA Jonathan Salmon Lockheed Missiles & Space, 1111 Lockheed Way, 80-41/102, Sunnyvale,

CA 94089-3504, USA

David Santiago Stanford U, Physics Dept, Stanford, CA 94305-7140, USA Ruggero M. Santilli Inst for Basic Research, P.O. Box 1577, Palm Harbor, FL 34682, USA Misao Sasaki Kyoto U, Dept of Physics, Kitashirakawa, Oiwake-cho, Kyoto, 606-01, JAPAN Katsuhiko Sato U of Tokyo, Dept of Physics, Bunkyo-ku, Tokyo 113, JAPAN Humitaka Sato Kyoto U, Dept of Physics, Kyoto 606-01, JAPAN James Satterthwaite 318 Franklin Ave, Cheltenham, PA 19012-2028, USA Peter Saulson Syracuse U, Dept of Physics, 201 Physics Bldg, Syracuse, NY 13244, USA Rick Savage Caltech, LIGO Project, MS 102-33, Pasadena, CA 91125, USA Marcelo Schiffer U Estadual de Campinas, Dep de Matematica Aplicada, CP 6065,

13081-970 Campinas S.P., BRAZIL

Roland Schilling Max-Planck-Institut, Hans-Kopfermann-Str 1, D-85748 Garching,

GERMANY

Christoph Schmid ETH Zürich, Inst for Theor Physics, CH-8093 Zürich, SWITZERLAND Richard Schoen Stanford U, Math Dept, Stanford, CA 94305, USA Bernard F. Schutz U of Wales, College of Cardiff, Dept of Physics, P.O. Box 913, Cardiff

CF2 3YB, UK

Susan Scott Australian Natl U, Centre for Math and its Applications, Canberra ACT 0200,

AUSTRALIA

Ed Seidel NCSA, 605 Springfield Ave, Champaign, IL 61820, USA Domenico Seminara M.I.T., Theor Physics, 77 Massachusetts Ave, Cambridge, MA 02139,

USA

Michael Shao Jet Propulsion Lab, 4800 Oak Grove Drive, MS 306-388, Pasadena, CA 91109,

USA

Peter Shelus U of Texas, McDonald Obs, Austin, TX 78712-6016, USA Yuri Shil’nov Kharkov State U, Dept Theor Physics, Svobody Sq 4, Kharkov, 31007,

UKRAINE

Kazunari Shima Saitama Institue of Techmology, Laoratory of Physics, Okabe-machi,

Saitama 338, JAPAN

Tetsuya Shimizu U of Tokyo, Dept of Physics, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113, JAPAN Hisa-aki Shinkai Waseda U, Dept of Physics, Okubo 3-4-1, Shinjuku, Tokyo 169-50, JAPAN Jonathan Simon U of Maryland, Physics Dept, College Park, MD 20742, USA Kenneth Skeldon U of Glasgow, Dept of Physics, University Ave, Glasgow G12 8QA

SCOTLAND, UK

Joel Smoller U of Michigan, Dept of Math, Ann Arbor, MI, USA George Smoot U of California, Lawrence Berkeley Lab and Space Sciences Lab, MS 5-232,

Berkeley, CA 94720, USA

Paulo M. Sõ Ue Do Algarve, Unidade de Ciencias Exactas e Humanas, Campus de Gambelas, P-8000 Faro, PORTUGAL Jiro Soda Kyoto U, Dept of Fundamental Sciences, FIHS, Nihonmatsu Cho, Yoshida, Kyoto

606-01, JAPAN

Pierre Sokolsky U of Utah, Physics Dept, Salt Lake City, UT 84112, USA Norbert Solomonson Louisiana State U, Dept of Physics, Baton Rouge, LA 70803, USA Shang Song U of New Mexico, Dept of Physics, 332 Pennington-Harbourton Rd, Pennington,

NJ 08534, USA

Euro Spallucci U of Trieste, Dept Theor Physics, Strada Costiera, 11, I-34014 Trieste, ITALY William Sparks Space Telescope Science Inst, 3700 San Martin Drive, Baltimore, MD 21218,

USA

Robert Spero Caltech, MS 102-33, Ligo Project, Pasadena, CA 91125, USA Philippe Spindel Universite de Mons Hainaut, 20, Place au Parc, B-7000 Mons, BELGIUM Frednik Ståhl Umeå U, Dept of Math, Axtorpsvagen 5E, S-90337 Umeå, SWEDEN Alexei Starobinsky Landau Inst for Theor Physics, Russian Academy of Sciences, Kosygina

2, Moscow 117334, RUSSIA

Robin Stebbins U of Colorado, CB #440, Boulder, CO 80309-0440, USA Alan Steif Cambridge U, DAMTP, Silver St, Cambridge CB3 9EW, UK Luigi Stella Osservatorio di Brera, via Brera, 28, I-20121 Milano,, ITALY Greg Stephens U of Maryland, Dept of Physics, College Park, MD 20742, USA Thomas R. Stevenson U of Maryland, Dept of Physics, College Park MD 20412, USA Ewan Stewart Kyoto U, Dept of Physics, Kyoto 606, JAPAN William Stoeger U of Arizona, Vatican Obs Research Group, Steward Obs, Tucson, AZ

85721, USA

Cosimo Stornaiolo INFN-Napoli, Mostra d’Oltremare Pad 19, I-80125 Napoli, ITALY Andy Strominger U of California, Dept of Physics, Santa Barbara, CA 93106, USA C. W. Stubbs U of California, Center for Particle Astrophysics, Berkeley, CA 94720, USA Wai-Mo Suen Washington U, Dept of Physics, St. Louis, Missouri 63130, USA R. Sunyaev Landau Inst, Moscow, RUSSIA S. V. Sushkov Kazan State Pedagogical Inst, Dept of Geometry, Mezhlauk St 1, 420021

Kazan, Tatarstan, RUSSIA

Paul Swanson Jet Propulsion Lab, 4800 Oak Grove Drive, Pasadena, CA, USA Michael Taber Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Takashi Tachizawa Waseda U, Dept of Physics, Okubo 3-4-1, Shinjuku-ku, JAPAN, JAPAN Ryu Takahashi Inst of Space, and Astronautical Science, 3-1-1 Yoshinodai, Sagamihara,

Kanagawa 229, JAPAN

Masaaki Takahashi Aichi U, Dept of Physics, Hirosawa, Igaya-Tyo, Kariya 448, JAPAN Judy Tam Hong Kong Polytechnic, Dept of Applied Math, Hung Hom, Kawloon,

HONG KONG

Takahiro Tanaka Kyoto U, Dept of Physics, Sakyo-ku, Kyoto 606-01, JAPAN Tsunefumi Tanaka Montana State U, Dept of Physics, Bozeman, MT 59717, USA Masayuki Tanimoto Tokyo Inst of Technology, Dept of Physics, Oh-Okayama 2-12-1,

Meguro-ku, Tokyo, 152, JAPAN

Mark Tapley Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Abraham Taub U of California, Math Dept, Berkeley, CA 94720, USA Marco Tavani Princeton U, Physics Dept, Jadwin Hall, Princeton, NJ 08544, USA Brett Taylor Montana State U, Physics Dept, Bozeman, MT 59717, USA Dietmar Theiss U of Stuttgart, Keplerstrasse 11, D-70174 Stuttgart, GERMANY Massimo Tinto Jet Propulsion Lab, 4800 Oak Grove Drive, MS 161-241, Pasadena, CA

91109, USA

Roman Tomaschitz Hiroshima U, Dept of Physics, 1-3-1 Kagamiyama, Higashi-Hiroshima

724, JAPAN

Rodney Torii Stanford U, Hansen Labs, Stanford, CA 94305, USA Charles Torre Utah State U, Physics Dept, Logan, UT 84322-4415, USA Sergio Torres Lawrence Berkeley Lab, Dept of Astrophysics, Berkeley, CA 94720, USA Pierre Touboul ONERA, BP 72 29 Ave de la Div Leclerc, 92322 Chatillon, FRANCE Jennie Traschen U of Massachusetts, Physics Dept, Amherst, MA 01003, USA Aldo Treves SISSA, ISAS, Via Beirut, 2-4, I-34013 Trieste, ITALY Michael J. Trinkala Hudson Valley Comm College, RD 3 #7 Coons Road, Troy, NY

12180-9622, USA

Kimio Tsubono U of Tokyo, Dept of Physics, Bunkyo-ku, JAPAN Sachiko Tsuruta Montana State U, Physics Dept, Bozeman, MT 59717, USA Roh-Suan Tung Natl Central U, Dept of Physics, Chung-Li 32054, TAIWAN John Turneaure Stanford U, Hansen Labs, GP-B, Stanford, CA 94305-4085, USA Edwin L. Turner Princeton U, Dept of Astrophysics, Princeton, NJ 08544-1001, USA Slava G. Turyshev Jet Propulsion Lab, Caltech, M/S 301-230, 4800 Oak Grove Drive,

Pasadena, CA 91109, USA

Toshio Uchida Tohoku U, Astronomical Inst, Aramaki, Aoba-ku, Sendai 980-77, JAPAN Ken-ichi Ueda U of Electro-Comm, Inst for Laser Science, 1-5-1 Chofugaoka, Chofu, Tokyo

182, JAPAN

Claes Uggla Stockholm U, Dept of Physics, Box 6730, S-11385 Stockholm, SWEDEN John Uglum Stanford U, Physics Dept, Stanford, CA 94306-4060, USA Abraham Ungar North Dakota State U, Dept of Math, Fargo, ND 58105, USA Richard Van Patten Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Richard Vassar Lockheed, 0/92-05 B/253, 3251 Hanover St, Palo Alto, CA 94304, USA Thomas Vaughan U of Oklahoma, 440 W. Brooks, Room 131, Norman, OK 73071, USA Rafael Vera U de Concepcion, Dep Fisica, Fac Ciencias, Casilla 4009, Concepcion, CHILE Massimo Visco CNR, Ist di Fisica dello Spazio Interplanetario, Via Galileo Galilei, 27,

I-00044 Frascati, ITALY

Stefano Vitale U of Trento, Dept of Physics, I-38050 Povo (Trento), ITALY Wolfgang Vodel Tera U, Inst of Solid State Physics, Max-Wien-Platz 1, D-07743 Jena,

GERMANY

Stamatis Vokos U of Washington, Physics Dept FM-15, Seattle, WA 98195, USA Chris Vuille Embry-Riddle Aero Univ, Dept of Physics, Daytona Beach, FL 32114, USA Robert Wagoner Stanford U, Physics Dept, Stanford, CA 94305-4060, USA Yong-Jiu Wang Hunan Normal U, Dept of Physics, Changsha, Hunan, CHINA Suwen Wang Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA Henry Ward U of Glasgow, Dept of Physics, Glasgow G12 8QQ Scotland, UK Joseph Weber U of California, Physics Dept, Irvine, CA 92717, USA Thomas Weiler Vanderbilt U, Physics Dept, BOX 1807 STA B, Nashville, TN 37235, USA Rainer Weiss M.I.T., 18 Vassar St, Room 20B-145, Cambridge, MA 02139, USA Paul Wesson U of Waterloo, Dept of Physics, Waterloo, ON N2L 3G1, CANADA Norbert Wex Friedrich-Schiller-U, MPG-AG Gravitation, Max Wien-Platz 1, D-07743 Jena,

GERMANY

John Wheeler Princeton U, Physics Dept, Princeton, NJ, USA James T. Wheeler Utah State U, Physics Dept, Logan, UT 84322-4415, USA John Whelan U of California, Dept of Physics, Santa Barbara, CA 93106-9530, USA Janet Whitesell Montana State U, Dept of Physics, Bozeman, MT 59717, USA Bernard Whiting U of Florida, 94 Landells Road, Pascoe Vale 3044, AUSTRALIA Lee Ashton Wild U of Wales, Dept of Physics, College of Cardiff, P.O. Box 913, Cardiff

CF2 3YB, UK

Jeffrey G. Williams Brandon U, Dept of Math, Brandon, Manitoba R7A 6A9, CANADA Paul Williamson Lockheed Missiles & Space Co, 3251 Hanover St, Palo Alto, CA 94304, USA Jim Wilson Lawrence Livermore Natl Lab, L-35, Livermore, CA 94550, USA Louis Witten U of Cincinnati, Dept of Physics, Cincinnati, OH 45221-0011, USA Kent Wood Naval Research Lab, Code 7621 NRL, Washington, DC 20375, USA Robert Wood Trinity Western U, 7600 Glover Road, Langley, BC V3A 6H4, CANADA Eric Woolgar Dalhousie U, Dept of Math, Halifax, Nova Scotia B3H 3J5, CANADA Paul Worden Stanford U, Hansen Labs, Stanford, CA 94305, USA Yueming Xiao Stanford U, Hansen Labs, GP-B, Stanford, CA 94305, USA She Sheng Xue INFN-Milan, Via Celoria 16, I-20133 Milano, ITALY Ion-Alexis Yadigaroglu Stanford U, Physics Dept, Stanford, CA 94305, USA Norifumi Yamada Clarkson U, Physics Dept, Potsdam, NY 13699-5820, USA Kazuhiro Yamamoto Kyoto U, Dept of Physics, Sakyo-ku, Kyoto 606-01, JAPAN Jun’ichi Yokoyama Kyoto U, Yukawa Inst for Theor Physics, Gokanosho, Uji 611, JAPAN Ulvi Yurtsever Caltech, Jet Propulsion Lab, Theor Astrophysics 130-33, Pasadena, CA

91125, USA

Alexander Zakharov Kazan State U, Physics Dept, Lenin St 18, 420008 Kazan, Tatarstan,

RUSSIA

R. Zalaletdinov Inst of Nuclear Physics, Dept of Theor Physics, Ulugbek, 702132 Tashkent,

UZBEKISTAN CIS

Sergio Zerbini U of Trento, Dept of Physics, Via Sommarive,, I-38050 Povo (Trento), ITALY Ping Zhou Stanford U, Hansen Labs, GP-B, Stanford CA 94305, USA Winfried Zimdahl U Düsseldorf, Inst fur Theor Physik, D-40225 Düsseldorf, GERMANY Michael Zucker Caltech, Dept of Physics, 102-33, Pasadena, CA 91125, USA